# Composite Learning Backstepping Control With Guaranteed Exponential Stability and Robustness

### v3

@@ I Introduction
Adaptive control is desirable due to its unique capacity to accommodate uncertain and time-varying properties of nonlinear systems, where recent survey papers can be referred to [31, 6, 42, 3, 46, 41]. The presence of mismatched uncertainties is a major obstacle to adaptive control of nonlinear systems. Adaptive integral backstepping with overparameterization, which combines integral backstepping and direct adaptive control, is a precursor to relax the above obstacle by designing an adaptive law to adjust a virtual control input at each backstepping step [18]. A tuning function approach is a direct adaptive backstepping approach without overparameterization [84, Ch. 4], where an adaptive law termed as a tuning function is constructed iteratively at each backstepping step, while an actual adaptive law is generated at the last step by all previous tuning functions. It is revealed that adaptive backstepping control driven by tuning functions has a higher-order tracking property [32]. There exist two common drawbacks for the above adaptive backstepping approaches: 1) The “explosion of complexity” exists due to the repeated differentiation of virtual control inputs; 2) the exponential stability of the closed-loop system (implying parameter convergence and robustness) relies on a strict condition termed persistent excitation (PE), which requires system states contain sufficiently rich spectral information all the time.
@@
Adaptive dynamic surface control (DSC) applies a first-order linear filter to estimate the time derivative of each virtual control input at its backstepping step for addressing the complexity problem of the classical backstepping control [49]. The performance and robustness of adaptive DSC have been enhanced by integrating neural network (NN) approximation [70], prescribed performance control [4], nonlinear filtering [79], coordinate transformation [77], power integration [26], etc. An approach similar to DSC, named command-filtered backstepping control (CFBC), employs second-order linear filters to estimate the time derivatives of virtual control inputs and compensation terms for stability guarantees [14]. The performance and robustness of adaptive CFBC have been improved by incorporating NN approximation [15], exact differentiation [47], immersion and invariance [16], etc. Yat, adaptive DSC and CFBC suffer from the explosion of dynamic order and the loss of global stability and asymptotic tracking due to filtering operations [50].
@@
A modular backstepping approach follows the certainty equivalence principle that separates control and estimation designs [84, Ch. 5]. A key feature of this approach is that the time derivatives of virtual control inputs are replaced by their partial derivatives with respect to system states and reference signals, while the resulting high-order time derivatives of parameter estimates are treated as additive disturbances. Thus, the modular backstepping approach does not involve tuning functions or overparameterization and has lower complexity. This approach ensures closed-loop stability with strong robustness by introducing a nonlinear damping term in a stabilizing function at each backstepping step [71]. A standard gradient-descent identifier derived from a swapping scheme can be combined with modular backstepping to achieve asymptotic tracking [84, Ch. 6]. Nevertheless, as the high-order time derivatives of parameter estimates exist in the closed-loop system, the modular backstepping approach may degrade transient tracking and prevent exact parameter estimation even in the presence of the PE condition.
@@
A high-order tuner (HOT) approach can efficiently remove the negative influence caused by the time derivatives of parameter estimates in modular backstepping [45]. The key idea of the HOT is to apply a linear filter with a sufficiently high relative degree to the adaptive law, such that the exact implementation of the high-order time derivatives of parameter estimates becomes feasible. In [30], a direct adaptive control scheme was combined with the HOT to counteract the transient process of parameter estimates caused by their high-order time derivatives, improving the transient and steady-state tracking performance. In [43], a memory regressor extension (MRE) identifier, which utilizes regressor extension with filtering, was combined with the HOT to design an indirect adaptive control law, where the HOT is applied to an extended regression model such that the high-order time derivatives of parameter estimates can be calculated exactly without filtering delay. However, the above two methods still need nonlinear damping terms to ensure closed-loop stability and transient performance and resort to the stringent PE condition for exponential stability guarantees.
@@
    The stringent PE condition must be fulfilled to realize the exponential stability of the closed-loop system.
@@
Notations: $\min\{\cdot\}$ denote the minimum operator, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ are the minimal and maximum eigenvalues of $A$, respectively, $\sigma_{\min}(A)$ is the minimum singular value of $A$, $\|\bm{x}\|$ is the Euclidean norm of $\bm{x}$, $L_{\infty}$ is the space of bounded signals, $I$ is an identity matrix, $\bm{0}$ is a zero matrix, $\Omega_{c}$ :=:= $\{\bm{x}|\|\bm{x}\|\leq c\}$ is the ball of radius $c$, $\arg\max_{x\in S}f(x)$ :=:= $\{x\in S|f(y)$ $\leq$ $f(x)$, $\forall y$ $\in$ S}S\}, $\bm{g}\in\mathcal{C}^{k}$ indicates that $\bm{g}$ has continuous partial derivatives up to the order $k$, where $A\in\mathbb{R}^{n\times n}$, $\bm{x}\in\mathbb{R}^{n}$, $c\in\mathbb{R}^{+}$, $f:\mathbb{R}\mapsto\mathbb{R}$, $\bm{g}$ :: $\mathbb{R}^{n}\mapsto\mathbb{R}^{m},$ $S\subset\mathbb{R}$, $n,m$ are positive integers, and $k$ is an non-negative integer.
@@ III Modular Backstepping Control Design
The following assumptions presented in [84] are given for the modular backstepping control design.
@@
and define an output vector $\bm{p}(t)$ :=:= $\bm{e}(t)+\bm{\zeta}(t)\in\mathbb{R}^{n}$, where $\bm{\zeta}(t)\in\mathbb{R}^{n}$ and ${\Phi}_{\rm s}(t)\in\mathbb{R}^{N\times n}$ are filtered outputs. Following the swapping technique [84, Ch. 6] and (10), one can obtain a static linear parametric model as a regression equation as follows:
@@
Remark 2: In existing modular backstepping control methods, nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ must be incorporated into the virtual control inputs $v_{i}$ in (5) to counteract the transient response arising from the modeling error term $\Phi^{T}\tilde{\bm{\theta}}$ and the high-order time derivatives $\hat{\bm{\theta}}^{(k)}$ ($k=1$ to n1n-1), so the boundedness of the closed-loop system can always be established even in the absence of adaptation, where $k_{\rm{d}\it i}\in\mathbb{R}^{+}$ are damping parameters ($i=1$ to $n$) [84, 43]. However, the square terms $\|\bm{\psi}_{i}\|^{2}$ can lead to high-gain control, regardless of whether the estimation error $\tilde{\bm{\theta}}$ converges to $\bm{0}$ or not, which results in noise amplification and control saturation in practice. In this study, we aim to establish closed-loop stability without resorting to nonlinear damping terms as detailed in Sec. IV.
@@ IV-A Composite Learning High-Order Tuner
with $\alpha_{i}\in\mathbb{R}^{+}$ ($i=$ 1 to $n$) being filtering constants to (13), which results in a generalized parameterized model
@@
Since the system (4) is continuous-time and Algorithm 1 is discrete-time implemented with the sampling time $T_{\rm s}$, $\tau_{\rm d}$ should be chosen to be greater than $T_{\rm s}$ to ensure the correct functioning of Algorithm 1. In Algorithm 1, we first choose a sufficiently small threshold $\sigma$ [see Line 1 in Algorithm 1]. Note that partial IE usually exists; otherwise, all channels will be deactivated. At the beginning of each partial IE stage, the maximal exciting strength $\sigma_{\rm c}$ is reset to $\sigma$ [see Line 7 in Algorithm 1]. If the current exciting strength $\sigma_{\min}(\Psi_{\zeta}(t))$ is greater than $\sigma_{\rm c}$, $t_{\rm e}$ and $\sigma_{\rm c}$ are updated [see Lines 9–11 in Algorithm 1]; otherwise, they remain unchanged. Thus, Algorithm 1 ensures that the exciting strength $\sigma_{\min}(\Psi_{\zeta}(t_{\rm e}))$ is monotonically non-decreasing at each partial IE stage. Fig. 1 illustrates $\sigma_{\rm c}$ in Algorithm 1 for a simple case with two partial IE stages. As new active channels $\bm{\phi}_{{\rm s},k_{j}}(t)$ exist at $t=T_{\rm a}$, the sub-regressor $\Phi_{{\rm s},\zeta}$ with the excitation matrix $\Psi_{\zeta}(t)$ in (18) is reconstructed by all new active channels [see Lines 4–8 in Algorithm 1]. In this partial IE stage, as the exciting strength $\sigma_{\min}(\Psi_{\zeta}(t))$ can be time-varying [see the green dash line in Fig. 1], $t_{\rm e}$ is updated based on exciting strength maximization, i.e., $\max_{\tau\in[T_{\rm a},t]}\sigma_{\min}(\Psi_{\zeta}(\tau))$ [see Lines 9–11 in Algorithm 1 and the black solid line in Fig. 1]. This is the same for the IE stage [see Lines 13–15 in Algorithm 1 and the red area in Fig. 1].
@@
With the generalized prediction error $\bm{\xi}(t)$ in (19) and the prediction error $\bm{\epsilon}(t)$ in (22), design a composite learning HOT
$\dot{\hat{\bm{\theta}}}=\Gamma\left(\Phi_{\rm f}\bm{\epsilon}(t)+\kappa\bm{\xi}(t)\right)$ (23)
in which $\Gamma\in\mathbb{R}^{N\times N}$ is a positive-definite diagonal matrix of learning rates, and $\kappa\in\mathbb{R}^{+}$ is a weighting factor. Since $H(s)$ in (16) owns n1n-1 relative degrees, the time derivatives of $\hat{\bm{\theta}}$ in (23) up to the (n1n-1)th order can be implemented physically by a direct differentiation scheme [35]. More specifically, the high-order time derivatives of $\hat{\bm{\theta}}$ are calculated by
$\hat{\bm{\theta}}^{(k+1)}=\Gamma\left(\sum_{i=0}^{k}C_{k}^{i}s^{k-i}H(s)[\Phi]\bm{\epsilon}^{(i)}+\kappa\bm{\xi}^{(k)}\right)$ (24)
@@ V-A Parameter Convergence Results
Let $\Phi_{{\rm f},\zeta}$ :=:= $\bm{\phi}_{{\rm f},k_{1}},\bm{\phi}_{{\rm f},k_{2}},\cdots,\bm{\phi}_{{\rm f},k_{N_{\zeta}}}]^{T}$ $\in\mathbb{R}^{N_{\zeta}\times n}$ denote an active sub-regressor of $\Phi_{\rm f}$ in (21), where $\bm{\phi}_{{\rm f},k_{j}}$ is the $k_{j}$th channel. Define a parameter estimation error $\tilde{\bm{\theta}}_{\zeta}$ :=:= $\tilde{\theta}_{k_{1}},\tilde{\theta}_{k_{2}},\cdots$, $\tilde{\theta}_{k_{N_{\zeta}}}]^{T}$ $\in\mathbb{R}^{N_{\zeta}}$ regarding active channels. The following theorem shows parameter convergence results of this study.
Theorem 1: Let $0,t_{\rm f})$ with $t_{\rm f}\in\mathbb{R}^{+}$ be the maximal time interval for the existence of solutions of the system (4). For any given unknown parameter vector $\bm{\theta}\in\Omega_{{\rm c}_{\theta}}$ and $\Gamma=\Gamma^{T}>0$, the composite learning law of $\hat{\bm{\theta}}$ in (23) ensures:
@@ V-B Closed-Loop Stability Results
We introduce the following lemmas to facilitate the analysis of closed-loop stability, where Lemma 2 can be obtained by a simple extension of [13, Lemma 4].
@@
Theorem 2: For the system (4) under Assumptions 1–3 driven by the CLBC law (6) and (23) with $\bm{x}(0)\in\Omega_{{\rm c}_{0}}$ and $\hat{\bm{\theta}}(0)\in\Omega_{{\rm c}_{\theta}}$, there exist proper control parameters $k_{{\rm c}1}$ to $k_{{\rm c}n}$ in (5) and filtered parameters $\alpha_{1}$ to $\alpha_{n-1}$ in (16), such that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) has:
@@
    Uniform ultimate boundedness (UUB) on $t$ $\in$ $0,\infty)$;
@@
Remark 3: In existing modular backstepping control methods, the nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ ($i=1$ to $n$) can ensure the boundedness of the system (4) in the absence of adaptation. In the proposed CLBC in (6) with (23), the non-update of some elements of the parameter estimate $\hat{\bm{\theta}}$ in (23) implies that the corresponding channels are inactive. In this case, partial IE in Definition 3 exists, and partial exponential stability of the closed-loop system without resorting to $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ can be established under Algorithm 1. It is worth noting in Theorem 2 that if the condition “the index set $\mathcal{I}$ no longer changes on $t\in(T_{\rm a},\infty)$” is not met, partial exponential stability of the closed-loop system can still be established. Actually, the endless changes of some indices $k_{j}$ imply the establishment of partial PE.
Remark 4: The proposed CLBC in (6) with (23) has several connections and distinctions compared to the MRE-HOT in [43]. Similarly to the MRE-HOT, the swapping and the stable filter $H(s)$ in (16) are applied to generate the high-order time derivatives $\hat{\bm{\theta}}^{(k+1)}$ in (24) for counteracting the negative effect due to the transient of parameter estimation. The distinctions between the two controllers include:
@@
    The proposed CLBC introduces the filtered prediction error $\bm{\epsilon}$ in (22) to counteract the transient response caused by the modeling error $\Phi^{T}\tilde{\bm{\theta}}$ in (7) such that closed-loop stability is established without the nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ ($i=1$ to $n$), whereas the MRE-HOT resorts to $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ to counteract $\Phi^{T}\tilde{\bm{\theta}}$ so as to establish UUB;

### v4

# Composite Learning Backstepping Control With Provable Exponential Stability and Robustness

@@ I Introduction
Adaptive control is desirable due to its unique capacity to accommodate uncertain and time-varying properties of nonlinear systems, where recent survey papers can be referred to [1, 2, 3, 4, 5, 6]. Due to its inherent advantages, adaptive control has been widely applied to various complex systems, such as Euler-Lagrange systems [7, 8], unmanned aerial vehicle systems [9], Markovian systems [10], and multi-agent systems [11, 12]. However, the presence of mismatched uncertainties is a major obstacle to adaptive control of nonlinear systems. Adaptive integral backstepping with overparameterization, which combines integral backstepping and direct adaptive control, is a precursor to relax the above obstacle by designing an adaptive law to adjust a virtual control input at each backstepping step [13]. A tuning function approach is a direct adaptive backstepping approach without overparameterization [14, Ch. 4], where an adaptive law termed as a tuning function is constructed iteratively at each backstepping step, while an actual adaptive law is generated at the last step by all previous tuning functions. It is revealed that adaptive backstepping control driven by tuning functions has a higher-order tracking property [15]. There exist two common drawbacks for the above adaptive backstepping approaches: 1) The “explosion of complexity” exists due to the repeated differentiation of virtual control inputs; 2) the exponential stability of the closed-loop system (implying parameter convergence and robustness) relies on a strict condition termed persistent excitation (PE), which requires system states contain sufficiently rich spectral information all the time [16].
@@
Adaptive dynamic surface control (DSC) applies a first-order linear filter to estimate the time derivative of each virtual control input at its backstepping step for addressing the complexity problem of the classical backstepping control [17]. The performance and robustness of adaptive DSC have been enhanced by integrating neural network (NN) approximation [18, 19], prescribed performance control [20], nonlinear filtering [21], coordinate transformation [22], power integration [23], etc. An approach similar to DSC, named command-filtered backstepping control (CFBC), employs second-order linear filters to estimate the time derivatives of virtual control inputs and compensation terms for stability guarantees [24]. The performance and robustness of adaptive CFBC have been improved by incorporating NN approximation [25], exact differentiation [26], immersion and invariance [27], etc. Yat, adaptive DSC and CFBC suffer from the explosion of dynamic order and the loss of global stability and asymptotic tracking due to filtering operations [28].
@@
A modular backstepping approach follows the certainty equivalence principle that separates control and estimation designs [14, Ch. 5]. A key feature of this approach is that the time derivatives of virtual control inputs are replaced by their partial derivatives with respect to system states and reference signals, while the resulting high-order time derivatives of parameter estimates are treated as additive disturbances. Thus, the modular backstepping approach does not involve tuning functions or overparameterization and has lower complexity. This approach ensures closed-loop stability with strong robustness by introducing a nonlinear damping term in a stabilizing function at each backstepping step [29]. A standard gradient-descent identifier derived from a swapping scheme can be combined with modular backstepping to achieve asymptotic tracking [14, Ch. 6]. Nevertheless, as the high-order time derivatives of parameter estimates exist in the closed-loop system, the modular backstepping approach may degrade transient tracking and prevent exact parameter estimation even in the presence of the PE condition.
@@
A high-order tuner (HOT) approach can efficiently remove the negative influence caused by the time derivatives of parameter estimates in modular backstepping [30]. The key idea of the HOT is to apply a linear filter with a sufficiently high relative degree to the adaptive law, such that the exact implementation of the high-order time derivatives of parameter estimates becomes feasible. In [31], a direct adaptive control scheme was combined with the HOT to counteract the transient process of parameter estimates caused by their high-order time derivatives, improving the transient and steady-state tracking performance. In [32], a memory regressor extension (MRE) identifier, which utilizes regressor extension with filtering, was combined with the HOT to design an indirect adaptive control law, where the HOT is applied to an extended regression model such that the high-order time derivatives of parameter estimates can be calculated exactly without filtering delay. However, the above two methods still need nonlinear damping terms to ensure closed-loop stability and transient performance and resort to the stringent PE condition for exponential stability guarantees. Online data memory provides a feasible approach to relax the excitation condition and improve parameter convergence and stability [33, 34], but existing results are based on the assumption that all regressor channels are activated in an uncorrelated manner.
@@
    The stringent excitation conditions must be fulfilled to realize the exponential stability of the closed-loop system.
@@
Notations: $\min\{\cdot\}$ denote the minimum operator, $\sigma_{\min}(A)$ is the minimum singular value of $A$, $\|\bm{x}\|$ is the Euclidean norm of $\bm{x}$, $L_{\infty}$ is the space of bounded signals, $I$ is an identity matrix, $\bm{0}$ is a zero matrix, $\Omega_{c}$ :=:= $\{\bm{x}|\|\bm{x}\|\leq c\}$ is the ball of radius $c$, $\arg\max_{x\in S}f(x)$ :=:= $\{x\in S|f(y)$ $\leq$ $f(x)$, $\forall y$ $\in$ S}S\}, $\bm{g}\in\mathcal{C}^{k}$ indicates that $\bm{g}$ has continuous partial derivatives up to the order $k$, where $A\in\mathbb{R}^{n\times n}$, $\bm{x}\in\mathbb{R}^{n}$, $c\in\mathbb{R}^{+}$, $f:\mathbb{R}\mapsto\mathbb{R}$, $\bm{g}$ :: $\mathbb{R}^{n}$ $\mapsto$ $\mathbb{R}^{m}$, $S\subset\mathbb{R}$, $n,m\in\mathbb{N}^{+}$, and $k\in\mathbb{N}$.
@@ III Modular Backstepping Control Design
The following assumption presented in [14] is given for the modular backstepping control design.
@@
and define an output vector $\bm{p}(t)$ :=:= $\bm{e}(t)+\bm{\zeta}(t)\in\mathbb{R}^{n}$, where $\bm{\zeta}(t)\in\mathbb{R}^{n}$ and ${\Phi}_{\rm s}(t)\in\mathbb{R}^{N\times n}$ are filtered outputs. Following the swapping technique [14, Ch. 6] and (10), one can obtain a static linear parametric model as a regression equation as follows:
@@
Remark 2: In existing modular backstepping control methods, nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ must be incorporated into the virtual control inputs $v_{i}$ in (5) to counteract the transient response arising from the modeling error term $\Phi^{T}\tilde{\bm{\theta}}$ and the high-order time derivatives $\hat{\bm{\theta}}^{(k)}$ ($k=1$ to n1n-1), so the boundedness of the closed-loop system can always be established even in the absence of adaptation, where $k_{\rm{d}\it i}\in\mathbb{R}^{+}$ are damping parameters ($i=1$ to $n$) [14, 32]. However, the square terms $\|\bm{\psi}_{i}\|^{2}$ can lead to high-gain control, regardless of whether the estimation error $\tilde{\bm{\theta}}$ converges to $\bm{0}$ or not, which results in noise amplification and control saturation in practice. In this study, we aim to establish closed-loop stability without resorting to nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ as detailed in Sec. IV.
@@ IV-A Composite Learning High-Order Tuner
with $\alpha_{i}\in\mathbb{R}^{+}$ ($i=$ 1 to n1n-1) being filtering constants to (13), which results in a generalized parameterized model
@@
Since the system (4) is continuous-time but Algorithm 1 is discrete-time implemented with the sampling time $T_{\rm s}$, $\tau_{\rm d}$ should be chosen to be greater than $T_{\rm s}$ to ensure the correct functioning of Algorithm 1. In Algorithm 1, we first choose a sufficiently small threshold $\sigma$ [see Line 1 in Algorithm 1]. Note that partial IE usually exists; otherwise, all channels will be deactivated. At the beginning of each partial IE stage, the maximal exciting strength $\sigma_{\rm c}$ is reset to $\sigma$ [see Line 7 in Algorithm 1]. If the current exciting strength $\sigma_{\min}(\Psi_{\zeta}(t))$ is greater than $\sigma_{\rm c}$, $t_{\rm e}$ and $\sigma_{\rm c}$ are updated [see Lines 9–11 in Algorithm 1]; otherwise, they remain unchanged. Thus, Algorithm 1 ensures that the exciting strength $\sigma_{\min}(\Psi_{\zeta}(t_{\rm e}))$ is monotonically non-decreasing at each partial IE stage. Fig. 1 illustrates $\sigma_{\rm c}$ in Algorithm 1 for a simple case with two partial IE stages. As new active channels $\bm{\phi}_{{\rm s},k_{j}}(t)$ exist at $t=T_{\rm a}$, the sub-regressor $\Phi_{{\rm s},\zeta}$ with the excitation matrix $\Psi_{\zeta}(t)$ in (18) is reconstructed by all new active channels [see Lines 4–8 in Algorithm 1]. In this partial IE stage, as the exciting strength $\sigma_{\min}(\Psi_{\zeta}(t))$ can be time-varying [see the green dash line in Fig. 1], $t_{\rm e}$ is updated based on exciting strength maximization, i.e., $\max_{\tau\in[T_{\rm a},t]}\sigma_{\min}(\Psi_{\zeta}(\tau))$ [see Lines 9–11 in Algorithm 1 and the black solid line in Fig. 1]. This is the same for the IE stage [see Lines 13–15 in Algorithm 1 and the red area in Fig. 1].
@@
With the generalized prediction error $\bm{\xi}$ in (19) and the filtered prediction error $\bm{\epsilon}$ in (22), design a composite learning HOT
$\dot{\hat{\bm{\theta}}}=\kappa_{1}\Phi_{\rm f}\bm{\epsilon}(t)+\kappa_{2}\bm{\xi}(t)$ (23)
in which $\kappa_{1}$, $\kappa_{2}\in\mathbb{R}^{+}$ are weighting factors. Since $H(s)$ in (16) owns n1n-1 relative degrees, the time derivatives of $\hat{\bm{\theta}}$ in (23) up to the (n1n-1)th order can be implemented physically by a direct differentiation scheme [38]. More specifically, the high-order time derivatives of $\hat{\bm{\theta}}$ are calculated by
$\hat{\bm{\theta}}^{(k+1)}=\kappa_{1}\sum_{i=0}^{k}C_{k}^{i}s^{k-i}H(s)[\Phi]\bm{\epsilon}^{(i)}+\kappa_{2}\bm{\xi}^{(k)}$ (24)
@@ V-A Parameter Convergence Results
Let $\Phi_{{\rm f},\zeta}$ :=:= $\bm{\phi}_{{\rm f},k_{1}},\bm{\phi}_{{\rm f},k_{2}},\cdots,\bm{\phi}_{{\rm f},k_{N_{\zeta}}}]^{T}$ $\in\mathbb{R}^{N_{\zeta}\times n}$ denote an active sub-regressor of $\Phi_{\rm f}$ in (21), in which $\bm{\phi}_{{\rm f},k_{j}}$ is the $k_{j}$th channel. Define a parameter estimation error $\tilde{\bm{\theta}}_{\zeta}$ :=:= $\tilde{\theta}_{k_{1}},\tilde{\theta}_{k_{2}},\cdots$, $\tilde{\theta}_{k_{N_{\zeta}}}]^{T}$ $\in$ $\mathbb{R}^{N_{\zeta}}$ regarding active channels. The following theorem shows parameter convergence results of this study.
Theorem 1: Let $0,t_{\rm f})$ with $t_{\rm f}\in\mathbb{R}^{+}$ be the maximal time interval for the existence of solutions of the system (4). For any given unknown parameter vector $\bm{\theta}\in\Omega_{{\rm c}_{\theta}}$, the composite learning law of $\hat{\bm{\theta}}$ in (23) ensures:
@@ V-B Closed-Loop Stability Results
We introduce the following lemmas to facilitate the analysis of closed-loop stability, where Lemma 2 can be obtained by a simple extension of [39, Lemma 4].
@@
Theorem 2: For the system (4) under Assumptions 1–2 driven by the CLBC law (6) and (23) with $\bm{x}(0)\in\Omega_{{\rm c}_{0}}$ and $\hat{\bm{\theta}}(0)\in\Omega_{{\rm c}_{\theta}}$, there exist proper control parameters $k_{{\rm c}1}$ to $k_{{\rm c}n}$ in (5) and filtered parameters $\alpha_{1}$ to $\alpha_{n-1}$ in (16), such that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) has:
@@
    Stability in the sense of uniform ultimate boundedness (UUB) on $t$ $\in$ $0,\infty)$;
@@
Remark 3: In existing modular backstepping control methods, the nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ ($i=1$ to $n$) can ensure the boundedness of the system (4) in the absence of adaptation. In the proposed CLBC in (6) with (23), the non-update of some elements for the parameter estimate $\hat{\bm{\theta}}$ in (23) implies that the corresponding channels are inactive. In this case, partial IE in Definition 3 exists, and partial exponential stability of the closed-loop system without resorting to $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ can be established under Algorithm 1. It is worth noting in Theorem 2 that if the condition “the index set $\mathcal{I}$ no longer changes on $t\in(T_{\rm a},\infty)$” is not met, partial exponential stability of the closed-loop system can still be established. Actually, the endless changes of some indices $k_{j}$ imply the establishment of partial PE.
Remark 4: The proposed CLBC in (6) with (23) has several connections and distinctions compared to the MRE-HOT in [32]. Similarly to the MRE-HOT, the swapping scheme and the stable filter $H(s)$ in (16) are applied to generate the high-order time derivatives $\hat{\bm{\theta}}^{(k+1)}$ in (24) for counteracting the negative effect due to the transient of parameter estimation. The distinctions between the two controllers include:
@@
    The proposed CLBC introduces the filtered prediction error $\bm{\epsilon}$ in (22) to counteract the transient response caused by the modeling error $\Phi^{T}\tilde{\bm{\theta}}$ in (7) such that closed-loop stability is established without the nonlinear damping terms $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ ($i=1$ to $n$), whereas the MRE-HOT resorts to $k_{\rm{d}\it i}\|\bm{\psi}_{i}\|^{2}$ to counteract $\Phi^{T}\tilde{\bm{\theta}}$ so as to establish UUB stability;